OpenStudy (naveenbhatia1312):

Find all solutions in the interval [0, 2π). (6 points) sin^2x - cos^2x = 0

OpenStudy (mathstudent55):

Treat this as a quadratic equation, a difference of squares. Factor it, then set each factor equal to zero and solve.

OpenStudy (naveenbhatia1312):

would the final answer be pi/4 or are there more or different solutions. @mathstudent55

OpenStudy (mathstudent55):

pi/4 is a correct solution, but there are more solutions in that interval.

OpenStudy (mathstudent55):

Did you factor the equation?

OpenStudy (naveenbhatia1312):

pi/4, 3pi/4, 5pi/4, and 7pi/4?

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):



OpenStudy (naveenbhatia1312):

Find the exact value by using a half-angle identity. cos(-pi/8)

OpenStudy (mathstudent55):

\(\sin^2 x - \cos^2 x = 0\) \((\sin x + \cos x)(\sin x - \cos x) = 0\) \(\sin x = -\cos x\) or \(\sin x = \cos x\) x = pi/4, 5pi/4 or x = 3pi/4, 7pi/4

OpenStudy (mathstudent55):

What is the half-angle identity for cos?

OpenStudy (naveenbhatia1312):

?

OpenStudy (mathstudent55):

Your second problem is asking you to use the half-angle identity to find the cos(-pi/8) You must start with the half-angle identity for cosine. What is it?

OpenStudy (mathstudent55):

\(\cos \dfrac{\theta}{2} = \pm \sqrt{\dfrac{1 + \cos \theta}{2} }\)

OpenStudy (naveenbhatia1312):

ohh i didnt know that

OpenStudy (mathstudent55):

pi/8 = (pi/4) / 2 Use theta = pi/4 since you know the exact value of cos pi/4 Substitute the values into the half-angle identity above. Then you need to choose the sign since the identity has \(\pm\) in it. To choose the sign, figure out what quadrant -pi/8 is in and what the sign of the cosine is in that quadrant.

OpenStudy (mathstudent55):

sorry, but gtg